# RsGCN: Rescaling Enhances Generalization of GCNs for Solving Scalable Traveling Salesman Problems

### v3

@@ 5.6 Ablation Study
To validate the effectiveness of the Rescaling Mechanism, we set the following ablation configurations: (1) Rs×2: rescaling adjacent nodes and subgraph edges, i.e., RsGCN; (2) Rs×1: only rescaling adjacent nodes; (3) Rs×0: no rescaling; (4) 5-NN: Re2Opt with the 5-nearest neighbors as candidates. Figure 6 reveals that the Rescaling Mechanism is crucial for improving GCNs’ generalization, where 5-NN acts as the borderline and the gap to 5-NN is presented. Without rescaling edges, Rs×1 shows inferior generalization compared to Rs×2. Even worse, Rs×0 fails to provide effective guidance for TSPs with 1K nodes and more. Due to its quadratic computational complexity, Rs×0 fails to complete inference for TSP-5K/10K within GPU memory. Figure 6 shows that the overall convergence speed follows Rs×2 > Rs×1 > 5-NN > Rs×0, validating the efficient guidance of RsGCN (the y-axis represents the average tour length across 16 instances for each scale, and the x-axis represents the runtime). Furthermore, ablation study on Re2Opt is presented in Appendix F.
Figure 4: Comparison of the number of learnable parameters among neural sovlers.
Figure 5: Comparison of zero-shot generalization.
Figure 6: Comparison of Re2Opt’s Convergence speed on large-scale TSPs.
To visually assess heatmaps’ quality, Figure 8 displays the consistent local regions of the ordered heatmaps generated by Rs×2/×1/×0 on two instances with 500 and 2K nodes, respectively. Some parts of the main diagonal in Rs×0’ heatmap do not obtain ground-truth heats, which easily lead to local optima. While for Rs×1, although the coverage on the main diagonal is enhanced compared to Rs×0, it suffers from the excessive dispersion of heats, which cannot efficiently guide the post-search. Overall, Rs×2’s heatmaps obtain the best coverage and the heat distribution is concentrated around the main diagonal, indicating its effectiveness and superiority.
Furthermore, two metrics are incorporated to evaluate heatmaps: (1) Average Top-k: The average rank of ground-truth nodes’ heats, and (2) Missing Rate in Top-5: The probability that ground-truth nodes are not included in the 5-hottest neighbors. Figure 8 presents the results of these two metrics, where Dist-1 denotes heatmaps obtained by calculating the reciprocal of distance matrices. Figure 8 indicates that Rs×2 achieves superior overall performance, further validating the effectiveness of the Rescaling Mechanism.
Figure 7: Comparison of ordered heatmaps.
@@ 6 Conclusion
To enhance the generalization capability of GCNs for scalable TSPs, we propose a new RsGCN that incorporates a Rescaling Mechanism.The Rescaling Mechanism includes rescaling adjacent nodes and rescaling subgraph edges to significantly reduce GCN’s sensitivity to scale-dependent features and help RsGCN learn universal patterns across TSPs of various scales. By 3-epoch training on a mixed-scale dataset composed of instances with up to 100 nodes, RsGCN can be successfully generalized to 10K-node instances without any fine-tuning. To efficiently utilize the heatmaps generated by RsGCN, a new post-search algorithm Re2Opt is proposed, which includes a reconstruction process to enhance search diversity and avoid local optima. Extensive experimental results on uniform distribution instances of 9 different scales from 20 to 10K and 78 real-world instances from TSPLIB validate the state-of-the-art performance of the proposed solver. In addition, the training cost and the number of learnable parameters are both significantly lower than those of the existing neural TSP solvers.
@@ Appendix A Ordered Heatmaps
Input: The original heatmap $\boldsymbol{H}$,
          the optimal tour $\hat{\Pi}$
@@ Appendix B Architecture of GCN
Our model adopts Residual Gated Graph Convnets [2, 12]. With regard to the global representation, both $i$ and $j$ index the node set $\mathcal{X}_{n}$ in this section.
@@
According to the problem definition, $x_{i}$ denotes the 2-dimensional unscaled coordinates in $\mathcal{X}_{n}$. Let $e_{i,j}$ denote the Euclidean distance between $x_{i}$ and $x_{j}$ ($\in\mathcal{N}_{i}^{k_{1}}$) based on the corresponding rescaled coordinates in $R(\mathcal{N}_{i}^{k_{1}})$. Subsequently, the unscaled coordinate $x_{i}$ and the rescaled edges $e_{i,j}$ are linearly embedded to $h$-dimensional features as follows:
@@
where $\boldsymbol{W}_{1}\in\mathbb{R}^{2\times h}$ and $\boldsymbol{W}_{2}\in\mathbb{R}^{1\times h}$ are learnable parameter matrices, $b_{1}$ and $b_{2}$ are learnable biases. Such a design of the embedding layer preserves global node distribution while ensuring stable feature representation of subgraphs during GCN’s message aggregation.
@@
where $\boldsymbol{W}_{3\sim 7}\in\mathbb{R}^{h\times h}$ are four learnable parameter matrices. LN denotes Layer Normalization, and GELU denotes the Gaussian Error Linear Unit used as the feature activation function. $\sigma$ denotes the Sigmoid function used as the gating activation function. $\odot$ denotes the Hadamard product operation.
@@
Let $e_{i,j}^{l}$ denote the edge features in the last layer, where $l$ is the maximum number of graph convolutional layers. $e_{i,j}^{l}$ is projected to $H_{i,j}\in[0,1]$ as follows:
$H_{i,j}=\begin{cases}\sigma(\text{MLP}(e_{i,j}^{l})),&\text{if }x_{j}\in\mathcal{N}_{i}^{k_{1}}\text{ and }i\neq j\\ 0,&\text{others}\end{cases}$ (11)
where MLP is a two-layer perceptron with GELU as the activation function. $H_{i,j}$ represents the heat between nodes $i$ and $j$, i.e., the probability that the edge connecting nodes $i$ and $j$ exists in the TSP tour.
@@ Appendix C Reproduction Details on Baselines
For neural solvers, we set the maximum batch size within the GPU memory limits to fully utilize the GPU. We set the maximum number of threads to 128 for CPU-dependent computation solvers Concorde, LKH-3, 2Opt, MCTS, and Re2Opt. For fairness, we adjust the hyperparameters to keep the runtime of baselines approximately the same. However, it remains difficult to ensure complete fairness, as the primary computational resources differ among solvers. LC solvers leverage the GPU’s powerful parallel processing capability to perform larger batch computations (typically covering all instances of each scale in one batch), whereas RC solvers mainly employ the GPU when generating heatmaps, with the post-search algorithms MCTS and Re2Opt relying entirely on the CPU.
@@
Among neural solvers, H-TSP, DRHG, DIFUSCO, and Fast T2T employ fine-tuning strategies. We conduct evaluation using the fine-tuned DRHG for TSP-1K+. H-TSP, DIFUSCO, and Fast T2T are all fine-tuned for multiple specific instance scales. For these three solvers, if no fine-tuned model corresponding to the test scale is available, we use model weights with the fine-tuned scale closest to the test scale for evaluation.
Appendix D Additional Results on TSP-20/50/100/200/500
Table 3: Comparison results on TSP-20/50/100/200/500.
Solver Type TSP-20 TSP-50 TSP-100 TSP-200 TSP-500
Concorde Exact 3.8403 0.0000 1.19s 5.6870 0.0000 3.35s 7.7560 0.0000 18.96s 10.7191 0.0000 7.44s 16.5464 0.0000 31.04s
LKH-3 Heuristic 3.8403 0.0000 28.32s 5.6870 0.0000 56.52s 7.7561 0.0003 55.33s 10.7193 0.0022 7.41s 16.5529 0.0391 10.12s
2Opt Heuristic 3.8459 0.1452 0.04s 5.7812 1.6490 0.07s 8.0124 3.3033 0.15s 11.2332 4.7946 0.08s 17.5688 6.1776 0.03s
H-TSP LC 10.8276 1.0122 6.09s 17.5962 6.3446 17.82s
LEHD LC 3.8403 0.0000 10.00s 5.6870 0.0005 25.00s 7.7572 0.0147 45.00s 10.7309 0.1103 15.00s 16.6603 0.6881 54.00s
GLOP LC 3.8406 0.0078 14.07s 5.6935 0.1143 29.42s 7.7839 0.3597 1.65m 10.7952 0.7099 19.86s 16.9089 2.1907 46.61s
DRHG LC 3.8403 0.0000 10.00s 5.6872 0.0039 25.00s 7.7581 0.0272 45.00s 10.7556 0.3410 15.00s 16.6448 0.5947 50.00s
Att-GCN RC+MCTS 3.8403 0.0000 2.39s + 10.14s 5.6875 0.0090 15.91s + 23.16s 7.7571 0.0137 24.21s + 46.13s 10.7625 0.4035 20.62s + 12.99s 16.7996 1.5308 31.17s + 51.53s
SoftDist MCTS 3.8745 0.8801 0.12s + 9.82s 5.7692 1.4495 0.12s + 23.93s 7.9553 2.5842 0.12s + 45.89s 10.8971 1.6647 0.12s + 13.39s 16.8161 1.6295 0.12s + 51.53s
DIFUSCO RC+MCTS 16.6378 0.5519 3.61m + 51.40s
Fast T2T RC+G+2Opt 3.8416 0.0348 53.00s 5.6876 0.0115 1.00m 7.7587 0.0337 1.23m 10.7445 0.2364 28.00s 16.6461 0.6024 37.00s
Re2Opt Heuristic 3.8403 0.0000 10.31s 5.6874 0.0066 22.40s 7.7626 0.0854 42.14s 10.7473 0.2634 12.46s 16.6309 0.5108 27.15s
[…]
Appendix E Additional Results on TSPLIB
@@
Tables 4 and 5 present the specific optimality gaps on 78 TSPLIB instances, where the results of DIFUSCO, T2T, and Fast T2T are sourced from [19], while those of POMO, BQ, LEHD, and DRHG are extracted from [17]. Due to the distinctive and unusual node distributions of instances fl1577 and fl3795, we activate all neighbors as candidate nodes in Re2Opt’ State Initialization only for solving fl1577 and fl3795. Consequently, the optimality gaps decrease as follows: For Re2Opt, 19.53% $\rightarrow$ 0.71% on fl1577 and 15.23% $\rightarrow$ 1.40% on fl3795; For RsGCN, 8.23% $\rightarrow$ 0.37% on fl1577 and 7.45% $\rightarrow$ 0.97% on fl3795. Note that without such a small adjustment on these two instances, the average optimality gap across all 78 instances obtained by the proposed RsGCN will grow to 0.49, which is still the best compared to other neural solvers.
@@
Instance DIFUSCO T2T Fast T2T POMO BQ LEHD DRHG Re2Opt RsGCN
@@
Instance DIFUSCO T2T Fast T2T POMO BQ LEHD DRHG Re2Opt RsGCN
@@
Appendix F Additional Ablation Study on Re2Opt
@@
Figure 10 shows the average optimal gaps (%) under different combinations of two hyperparameters: the maximum number of actions $M$ and candidate set sizes $k_{2}$. We select 16 instances for each scale. The runtime limit of Re2Opt is set to $0.05n$ seconds for $n$-node instances. The x-axis represents $M$ with different sampling ranges $[10,20]$, $[10,30]$, $[10,40]$, $[10,50]$, and $[10,60]$, while the y-axis represents $k_{2}$ with different values across $[4,7]$. As shown in Figure 10, $k_{2}$ has a more significant impact on Re2Opt. For larger instance scales, a smaller $k_{2}$ tends to help Re2Opt achieve smaller optimality gaps. As a larger $k_{2}$ expands the search space, the efficiency of reconstruction is weakened and the runtime of each 2Opt step is prolonged in dealing with large-scale instances.
Figure 10: Ablation study on the two hyperparameters of Re2Opt.
Appendix G Comparisons of Experimental Configurations
@@
Table 6 presents the configurations of the main training phase for each neural solver, where SL denotes supervised learning and RL denotes reinforcement learning. Our proposed RsGCN requires significantly fewer learnable parameters and training epochs compare to other baselines, substantially reducing the training cost. In GLOP, “×3” indicates that GLOP requires combining 3 neural models with identical parameter sizes during a single solving process, whereas the “+” in H-TSP denotes the combination of two neural models with different parameter scales. Both GLOP and H-TSP employ reinforcement learning, requiring a very large number of training epochs.
Table 6: Comparisons on training configurations for neural solvers.
Solver Learning Paradigm Learnable Parameters Training Epochs
RsGCN SL 0.417M 3
LEHD SL 1.43M 150
DRHG SL 2.65M 100
GLOP RL 1.30M × 3 >500
DIFUSCO SL 5.33M 50
Fast T2T SL 5.33M 50
H-TSP RL 5.34M + 2.51M >500
Att-GCN SL 11.1M 15
Appendix H Visualization of Rescaling Mechanism
Figure 11 visually illustrates the influence of scale-dependent features and the effect of the Rescaling Mechanism. In the normalized square unit space, the distribution of nodes shown in the first row becomes denser as the number of nodes increases. Since nodes are fully connected in TSPs, the number of adjacent nodes for each node is a scale-dependent feature. To rescale adjacent nodes, k-Nearest Neighbor Selection is conducted and each node forms a subgraph containing itself and its top five nearest neighbors in the instance. Note that Figure 11 presents the subgraph of a given node (denoted by orange color) for each instance in the second row. By rescaling adjacent nodes, the subgraphs of nodes from TSP instances with various scales have a uniform number of adjacent nodes, which achieves graph sparsification and helps RsGCN learn universal patterns to enhance generalization capability.
In addition, we observe that the magnitude of edge lengths in the subgraphs still varies in different scales of instances, thereby the length of subgraph edge is another scale-dependent feature. As shown in Figure 11, the average distances between the given node and its neighbors in the three instances are 0.1793, 0.0878, and 0.0206 respectively, which have a significant difference in magnitude and will weaken the generalization capability. Note that the corresponding positions of the subgraphs are denoted by red frames in the distribution of nodes shown in the first row, which intuitively reveal the magnitude difference. Thus, the Rescaling Mechanism further rescales the subgraph edges by Uniform Unit Square Projection to obtain the rescaled subgraph shown in the third row. We can see that the average distances between the given node and its neighbors are in the same magnitude for the three instances, while the topological structure of the subgraphs is preserved. By rescaling subgraph edges, the lengths of subgraph edges across TSPs of various scales are adjusted to the same magnitude, which helps RsGCN generalize the patterns learned from small-scale instances for large-scale instances.
Figure 11: Visualization of the Rescaling Mechanism on TSP-20/200/2000.
Appendix I Limitations and Broader Impacts

### v4

@@ 5.6 Ablation Study
To validate the effectiveness of the subgraph-based rescaling, we set the following ablation variants: (1) Rs×2: applying subgraph-based rescaling, i.e., RsGCN; (2) Rs×1: constructing subgraphs without rescaling edge lengths; (3) Rs×0: disabling subgraph-based rescaling, equivalent to the vanilla Gated GCN; (4) 5-NN: RBS with the 5-nearest neighbors as candidates. Figure 5 presents the gaps of each variant with respect to 5-NN, revealing that the subgraph-based rescaling is crucial for improving GCNs’ generalization. Without rescaling edges, Rs×1 shows inferior generalization compared to Rs×2. Even worse, Rs×0 fails to provide effective guidance for TSPs with more than 500 nodes. Figure 6 shows that the overall convergence speed follows Rs×2 > Rs×1 > 5-NN > Rs×0, validating the effective guidance of RsGCN (the y-axis represents the average tour length across 16 instances for each scale, and the x-axis represents the runtime). Furthermore, additional ablation study on RBS is presented in Appendix E.
Figure 4: Comparison of neural model sizes.
Figure 5: Comparison of guidance effect among ablation variants.
Two metrics are incorporated to evaluate heatmaps on unseen-scale TSPs: (1) Average Rank: The average rank of optimal successors’ heats, and (2) Missing Rate in Top-5: The probability that optimal successors are not included in the 5-hottest neighbors. Figure 7 presents the results of these two metrics, where Dist-1 denotes heatmaps obtained by calculating the reciprocal of distance matrices. Figure 7 indicates that Rs×2 achieves superior overall performance, further validating the effectiveness of the subgraph-based rescaling. Furthermore, to visually assess heatmaps’ quality, we introduce ordered heatmaps that are sorted according to the optimal tour. The detailed definition of the ordered heatmaps, along with the ablation results, is provided in Appendix A.
5.7 Computational Cost
Using publicly released data from previous studies, we compare the training cost of our RsGCN with Fast T2T (Li et al., 2024), as presented in Table 4. Fast T2T employs a Diffusion model as its decoder, which incurs additional training overhead. Moreover, it does not perform edge normalization, resulting in poor cross-scale generalization and requiring additional fine-tuning to generalize to large-scale TSPs. In contrast, our RsGCN achieves strong generalization while remaining lightweight, significantly reducing the training cost. Furthermore, Table 4 reports the single-batch inference time and memory usage of RsGCN across different problem scales. RsGCN exhibits good linear memory scalability, suggesting its potential to address ultra-large-scale TSPs. In addition, more comparisons of training configurations are presented in Appendix F.
Figure 6: Comparison of boost to RBS’s Convergence speed on TSP-1K/2K.
@@ 5.7 Computational Cost
Table 3: Comparison of training cost between Fast T2T and RsGCN.
Model Problem Scale Dataset Size Batch Size GPU Memory Time
Fast T2T TSP-50 1502K 32 1 × A100 16.5GB 112h 45m
TSP-100 1502K 32 23.2GB 488h 12m
RsGCN
TSP-Mix
{20,30,50,100}
1000K 32 1 × A100 4.0GB 50m
1 × H20 3.4GB 28m
Table 4: Time and memory usage during inference.
TSP- 50 100 500 1K 5K 10K 50K 100K 500K
Time (s) 0.01 0.01 0.01 0.01 0.01 0.01 0.11 0.32 2.01
Memory (GB) 0.02 0.02 0.08 0.15 0.74 1.46 7.28 14.55 72.70
@@ 6 Conclusion
To enhance the cross-scale generalization capability of GCNs for solving TSPs, we propose a new RsGCN that incorporates the subgraph-based rescaling. The subgraph-based rescaling includes constructing subgraphs and rescaling subgraph edges to help RsGCN learn universal patterns across various scales. By 3-epoch training on a mixed-scale dataset composed of instances with up to 100 nodes, RsGCN can generalize to 10K-node instances without any fine-tuning. Extensive experimental results on uniform-distribution instances of 9 different scales from 20 to 10K and 78 real-world instances from TSPLIB validate the state-of-the-art performance of our method. In addition, the training cost and the number of learnable parameters are both significantly lower than those of the neural baselines.
@@ Appendix A Ordered Heatmaps
Figure 9 displays the identical 100×100 local regions of the ordered heatmaps generated by Rs×2/×1/×0 on two instances of TSP-500 and TSP-2K, respectively. Some cells close to the main diagonal in Rs×0’ heatmap do not obtain heats, which easily lead to local optima. While for Rs×1, although the coverage around the main diagonal is enhanced compared to Rs×0, it suffers from the excessive dispersion of heats, which provides only limited guidance for the post-search. Overall, Rs×2’s heatmaps reveal a concentration of heat distribution around the main diagonal, demonstrating its effectiveness and superiority.
Input: The original heatmap $\boldsymbol{H}$, the optimal tour $\hat{\Pi}$
@@ Appendix B Architecture of GCN
Our model adopts Residual Gated Graph Convnets (Bresson & Laurent, 2018; Joshi et al., 2019).
@@
According to the problem definition, $x_{i}$ denotes the 2-dimensional coordinates in $\mathcal{X}_{n}$. Let $e_{i,j}$ denote the rescaled distance between $x_{i}$ and $x_{j}$ in subgraph $\mathcal{N}_{i}^{k_{1}}$. Subsequently, the global coordinate $x_{i}$ and the subgraph edge $e_{i,j}$ are linearly embedded to $h$-dimensional features as follows:
@@
where $\boldsymbol{W}_{1}\in\mathbb{R}^{2\times h}$ and $\boldsymbol{W}_{2}\in\mathbb{R}^{1\times h}$ are learnable parameter matrices, $b_{1}$ and $b_{2}$ are learnable biases. Such an embedding design retains the global distribution view while maintaining a unified subgraph perspective during GCN message aggregation.
@@
where $\boldsymbol{W}_{3\sim 7}\in\mathbb{R}^{h\times h}$ are four learnable parameter matrices. LN denotes Layer Normalization (Ba et al., 2016), and GELU denotes the Gaussian Error Linear Unit (Hendrycks & Gimpel, 2016) used as the feature activation function. $\sigma$ denotes the sigmoid used as the gating activation function. $\odot$ denotes the Hadamard product operation.
@@
Let $e_{i,j}^{l}$ denote the edge features in the last layer, where $l$ is the maximum number of graph convolutional layers. $e_{i,j}^{l}$ is projected to $\boldsymbol{H}_{i,j}\in(0,1)$ as follows:
$\boldsymbol{H}_{i,j}=\begin{cases}\sigma(\text{MLP}(e_{i,j}^{l})),&\text{if }x_{j}\in\mathcal{N}_{i}^{k_{1}}\text{ and }i\neq j\\ 0,&\text{others}\end{cases}$ (11)
where MLP is a two-layer perceptron with GELU as the activation function. $\boldsymbol{H}_{i,j}$ represents the probability that edge $x_{i}\rightarrow x_{j}$ exists in the predicted TSP tour.
@@ Appendix C Reproduction Details on Baselines
For neural methods, we set the maximum batch size within the GPU memory limits to fully utilize the GPU. We set the maximum number of threads to 128 for CPU-dependent computation solvers Concorde, LKH-3, 2-Opt, MCTS, and RBS. For fairness, we adjust the hyperparameters to keep the runtime of baselines approximately the same. However, it remains difficult to ensure complete fairness, as the primary computational resources differ among solvers. Transformer-based methods leverage the GPU’s powerful parallel processing capability to perform larger batch inference (typically covering all instances of each scale in one batch), whereas GCN-based methods mainly employ the GPU when generating heatmaps, with the post-search algorithms MCTS and RBS relying entirely on the CPU.
@@
Among neural methods, H-TSP, DRHG, DIFUSCO, and Fast T2T employ fine-tuning strategies. We conduct evaluation using the fine-tuned DRHG for TSP-1K+. H-TSP, DIFUSCO, and Fast T2T are all fine-tuned for multiple specific instance scales. For these three solvers, if no fine-tuned model corresponding to the test scale is available, we use model weights with the fine-tuned scale closest to the test scale for evaluation.
Appendix D Additional Results on TSPLIB
Note that we increase $k_{2}$ to 10 for TSPLIB, because RsGCN is trained only on uniform-distribution TSPs, making cross-distribution generalization to real-world TSPs challenging. We conduct an ablation study on the effect of $k_{2}$ on solving TSPLIB instances, reporting the minimum and average gaps over 5 repeated trials. As shown in Table 5, for the minimum gaps, the differences among $k_{2}\in\{5,7,10\}$ are minor, while for the average gaps, $k_{2}=10$ performs better. Overall, for instances from distributions that RsGCN has not been trained on, appropriately enlarging the RBS candidate set improves search robustness.
Table 5: The effect of $k_{2}$ on solving TSPLIB instances.
Size Min Gap(%) Avg Gap(%)
$k_{2}$=5 $k_{2}$=7 $k_{2}$=10 $k_{2}$=5 $k_{2}$=7 $k_{2}$=10
<<100 0.00 0.00 0.00 0.00 0.00 0.00
[100, 200) 0.01 0.00 0.00 0.24 0.23 0.23
[200, 500) 0.21 0.03 0.04 1.08 0.67 0.21
[500, 1K) 0.11 0.07 0.13 0.73 0.44 0.26
$\geq$1K 0.64 0.69 0.77 1.30 1.20 1.13
All 0.29 0.27 0.30 0.83 0.68 0.55
@@ Appendix D Additional Results on TSPLIB
Tables 6 and 7 present the specific optimality gaps on 78 TSPLIB instances, where the results of DIFUSCO, T2T, and Fast T2T are sourced from Li et al. (2024), while those of POMO, BQ, LEHD, and DRHG are extracted from Li et al. (2025). Due to the distinctive and unusual node distributions of instances fl1577 and fl3795, we activate all neighbors as candidate nodes in RBS’ State Initialization only for solving fl1577 and fl3795. Consequently, the optimality gaps decrease as follows: For RBS, 19.53% $\rightarrow$ 0.71% on fl1577 and 15.23% $\rightarrow$ 1.40% on fl3795; For RsGCN, 8.23% $\rightarrow$ 0.37% on fl1577 and 7.45% $\rightarrow$ 0.97% on fl3795. Note that without such a small adjustment on these two instances, the average optimality gap across all 78 instances obtained by the proposed RsGCN will grow to 0.49, which is still the best compared to other neural methods.
@@
Instance DIFUSCO T2T Fast T2T POMO BQ LEHD DRHG RBS RsGCN
@@
Instance DIFUSCO T2T Fast T2T POMO BQ LEHD DRHG RBS RsGCN
@@
Appendix E Additional Ablation Study on RBS
As described in Section 4.3.3, edge enhancement guides RBS toward more effective reconstructions, allowing it to overcome local optima. This strategy, which utilizes historical experience, has been used in both Monte Carlo search trees and ant colony algorithms. To validate the effect of edge enhancement, we conduct repeated experiments on large-scale TSPs with it turned off. As shown in Table 8, edge enhancement significantly improves the efficiency of RBS.
Table 8: The effect of edge enhancement in RBS on large-scale TSPs, where E and w/o E denote with and without edge enhancement, respectively.
Type TSP-2K TSP-5K TSP-10K
Length Gap(%) Length Gap(%) Length Gap(%)
E 32.6784 0.5698 51.3987 0.6645 72.6335 0.9103
w/o E 32.7679 0.8455 51.5379 0.9370 72.7828 1.1178
@@ Appendix E Additional Ablation Study on RBS
Figure 10 shows the average optimal gaps (%) under different combinations of two hyperparameters: the maximum number of actions $M$ and candidate set sizes $k_{2}$. We select 16 instances for each scale. The runtime limit of RBS is set to $0.05n$ seconds for $n$-node instances. The x-axis represents $M$ with different sampling ranges $[10,20]$, $[10,30]$, $[10,40]$, $[10,50]$, and $[10,60]$, while the y-axis represents $k_{2}$ with different values across $[4,7]$. As shown in Figure 10, $k_{2}$ has a more significant impact on RBS. For larger instance scales, a smaller $k_{2}$ tends to help RBS achieve smaller optimality gaps. As a larger $k_{2}$ expands the search space, the efficiency of reconstruction is weakened and the runtime of each 2-Opt step is prolonged in dealing with large-scale instances.
Figure 10: Ablation study on the hyperparameters $M$ and $k_{2}$ of RBS.
Appendix F Comparison of Training Configurations
@@
Table 9 presents the configurations of the main training phase for each neural solver, where SL denotes supervised learning and RL denotes reinforcement learning. Our RsGCN requires significantly fewer learnable parameters and training epochs compared to other baselines, substantially reducing the training cost. Moreover, our main results in Section 5.5 also demonstrate that supervised learning on small-scale TSPs is an effective training strategy for RsGCN. In GLOP, “×3” indicates that GLOP requires combining 3 neural models with identical parameter sizes during a single solving process, whereas the “+” in H-TSP denotes the combination of two neural models with different parameter scales. Both GLOP and H-TSP employ reinforcement learning, requiring a very large number of training epochs.
Table 9: Comparison on training configurations of neural methods.
Paradigm Solver #Parameters Training Epochs
SL RsGCN 0.417M 3
LEHD 1.43M 150
DRHG 2.65M 100
DIFUSCO 5.33M 50
Fast T2T 5.33M 50
Att-GCN 11.1M 15
RL GLOP 1.30M × 3 >500
H-TSP 5.34M + 2.51M >500
Appendix G Visualization of Subgraph-Based Rescaling
Figure 11 visually illustrates the effect of the subgraph-based rescaling. As shown in the first row, the node distributions become denser in the unit square as the number of nodes increases. First, 5-NN selection is conducted, and each node forms a subgraph containing itself and its 5-nearest neighbors. Then, we observe that the magnitude of edge lengths in subgraphs varies across different scales. As shown in the second row, the average distances between the given node and its neighbors in the three instances are 0.1793, 0.0878, and 0.0206 respectively. Thus, we employ Uniform Unit Square Projection to rescale subgraph edges, obtaining the rescaled subgraph shown in the third row. We can see that the average distances are rescaled to the same magnitude across different scales, while the proportion of the subgraphs is preserved. Overall, the subgraph-based rescaling enhances the cross-scale generalization of RsGCN.
Appendix H Limitations and Broader Impacts